Hello I need help sovlving the below problems: I'm not very good at word problems. Could you please help. I need to show my work also. Thank you
1) The sum of two numbers is greater than 12. The second number is 8 more than the first. What are the possible values for the first of the two numbers? In your answer, denote the first number by X.
2) Max purchased a prepaid phone card for $15. Long distance calls cost 7 cents a minute using this card. Mai used her card only once to make a long distance call. If the remaining credit on her card is $12.83, now many minutes did her call last?
3) Ed will rent a car for the weekend. He can choose one of two payment plans. The first plan costs $55.98 for 2 days plus 13 cents per mile. The second plan costs $63.98 for two days plus 11 cents per mile. How many miles does Ed need to drive for the two plans to cost the same?
4) A local hamburger shop sold a combined total of 802 hamburgers and cheeseburgers on Monday. There were 52 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday?

5) Solve for Z: 4z + 13 = 85

Answer by mananth(16949)

1) The sum of two numbers is greater than 12. The second number is 8 more than the first. What are the possible values for the first of the two numbers? In your answer, denote the first number by X.
let the number be x
the other number will be x+8
x+x+8>12
2x+8>12
2x>12-8
x>4/2
x>2
the other number is greater than 10
..
..
2) Max purchased a prepaid phone card for $15. Long distance calls cost 7 cents a minute using this card. Mai used her card only once to make a long distance call. If the remaining credit on her card is $12.83, now many minutes did her call last?
let the number of minutes used be x
7 cents /minute = $0.07
usage = 0.07*x
balance = 12.83
0.07x+12.83=15
0.07x=15-12.83
0.07x=2.17
x=31 minutes
..
..
3) Ed will rent a car for the weekend. He can choose one of two payment plans. The first plan costs $55.98 for 2 days plus 13 cents per mile. The second plan costs $63.98 for two days plus 11 cents per mile. How many miles does Ed need to drive for the two plans to cost the same?
let the number of miles he drives be x
first plan
55.98+0.13x
second plan
63.98+0.11x
..
55.98+0.13x=63.98+0.11x
0.02x=8
x=8/0.02
x=400 miles


4) A local hamburger shop sold a combined total of 802 hamburgers and cheeseburgers on Monday. There were 52 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Monday?
let x = hamburgers
x+52 = cheeseburgers
x+x+52=802
2x+52=802
2x=750
x=325 hamburgers
802-325= 477 cheeseburgers
5) Solve for Z: 4z + 13 = 85
4z+13-13=85-13
4z=72
z=18

M Ananth for regular help.